OpenStudy (anonymous):

Find the value of x. RT = 4x+10 SW = 5x-20

OpenStudy (anonymous):

What are RT and SW?

OpenStudy (anonymous):

just the names of the lines , its nothinq imortant only letter important is x

OpenStudy (anonymous):

r they diagonals of a rectangle/square or ????

OpenStudy (anonymous):

rectangle

OpenStudy (anonymous):

see diagonals of rectangle are equal therefore RT = SW 4x + 10 = 5x - 20

OpenStudy (anonymous):

further 10 + 20 = 5x - 4x 30 = x

OpenStudy (anonymous):

ok how about RT = 2x+4 Sw = 6x-20

OpenStudy (anonymous):

see if they are again diagonals of a rectangle then RT = SW equate them as i had done and solve to get value of x, it shud b easy since i have already done one....

OpenStudy (anonymous):

is it 6

OpenStudy (anonymous):

2x + 4 = 6x - 20 4 + 24 = 6x - 2x 28 = 4x 28/4 = x 7 = x

OpenStudy (anonymous):

pls click GOOD ANSWER button for me if u r satisfied with the answer... (: }))

OpenStudy (anonymous):

abcd is a rectangle . find the length od the diagona. ac=2x-3 bd=x+5

OpenStudy (anonymous):

-30

OpenStudy (anonymous):

length can never be negative !!!!!!!!!!